# Bridging Classical and Learning-based Iterative Registration through Deep Equilibrium Models

### v1

@@ 1 Introduction
  • We propose DEQReg, the first end-to-end unsupervised registration framework that formulates DIR as a fixed-point equilibrium problem. Extensive experiments on the public brain MRI and lung CT datasets demonstrate that our approach achieves competitive accuracy with $O(1)$ memory complexity through implicit differentiation.
@@ 2.2 Deep Equilibrium Models for Deformable Image Registration
where $I$ is an identity matrix of the same size as $\frac{\partial g_{\theta}}{\partial{\phi}^{*}}$. Leveraging IFT, equilibrium points can be derived using any black-box root-finding algorithm, e.g. Broyden’s method [7]. Fig. 2 provides a schematic comparison between DEQ and existing unrolling methods. By focusing on the equilibrium state rather than the iteration trajectory, DEQ achieves $O(1)$ memory complexity while enabling theoretically unlimited iterations, with convergence design (i.e. fixed point condition).
@@ 5 Conclusion
For deformable medical image registration, classical optimization methods typically converge, but modern learning-based unrolling methods lack the theoretical guarantee of convergence and suffer from memory bottlenecks. This creates a gap between the principled convergence of classical methods and the lack of convergence in learning-based methods. In this work, we propose DEQReg to bridge this gap by reformulating learning-based registration as an equilibrium-seeking problem. Our experiments show that this reformulation enables stable convergence, in contrast to existing unrolling methods that tend to degrade when further unrolled beyond the training steps. DEQReg achieves competitive accuracy and improved regularity with $O(1)$ memory usage, effectively unifying classical optimization principles and learning-based advantages in speed and flexibility.

### v2

@@ 1 Introduction
  • We propose DEQReg, the first end-to-end unsupervised registration framework that formulates DIR as a fixed-point equilibrium problem. Extensive experiments on the public brain MRI and lung CT datasets demonstrate that our approach achieves competitive accuracy with constant memory complexity through implicit differentiation.
@@ 2.2 Deep Equilibrium Models for Deformable Image Registration
where $I$ is an identity matrix of the same size as $\frac{\partial g_{\theta}}{\partial{\phi}^{*}}$. Leveraging IFT, equilibrium points can be derived using any black-box root-finding algorithm, e.g. Broyden’s method [7]. Fig. 2 provides a schematic comparison between DEQ and existing unrolling methods. By focusing on the equilibrium state rather than the iteration trajectory, DEQ achieves constant memory complexity while enabling theoretically unlimited iterations, with convergence design (i.e. fixed point condition).
@@ 5 Conclusion
For deformable medical image registration, classical optimization methods typically converge, but modern learning-based unrolling methods lack the theoretical guarantee of convergence and suffer from memory bottlenecks. This creates a gap between the principled convergence of classical methods and the lack of convergence in learning-based methods. In this work, we propose DEQReg to bridge this gap by reformulating learning-based registration as an equilibrium-seeking problem. Our experiments show that this reformulation enables stable convergence, in contrast to existing unrolling methods that tend to degrade when further unrolled beyond the training steps. DEQReg achieves competitive accuracy and improved regularity with constant memory usage, effectively unifying classical optimization principles and learning-based advantages in speed and flexibility.